OpenStudy (anonymous):

help please 8th grade math! Which statement is true? A. –10 + (–9) > 3 + (–25) B. –18 + 21 = –15 + 12 C. 13 + (–7) > 21 + (–14) D. 7 + (–6) < –14 + 12

OpenStudy (doggy):

B

OpenStudy (doggy):

because when you subtract a negative number with a positive number, you always use the sign of the bigger number

OpenStudy (anonymous):

Addition of a negative is the same as subtracting positive numbers. Using this, figure out the values of each side, then find the correct statement.

OpenStudy (anonymous):

OH okay thank you soo much!! Do you mind helping me with this one? What is the value of the variable that makes the statement true? b + (–2) + 8 = 0 A. –10 B. –6 C. 6 D. 10

OpenStudy (anonymous):

Remember what the last post said about adding negatives. Using that rule, find the value of the constant, then move it to the right side.

OpenStudy (anonymous):

huh?

OpenStudy (anonymous):

What would -6+-2 be?

OpenStudy (anonymous):

b is a variable, but (-2) and 8 are constants. Since we're adding a negative, we have to subtract 2 from 8. 8 - 2 = 6. Move the six over to the right side of the original equation, then find the value of b

OpenStudy (anonymous):

so would it be B?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

okay thank you so much!!

OpenStudy (anonymous):

Do you need help with any more?

OpenStudy (anonymous):

b

OpenStudy (anonymous):

yes one more. if you dont mind. :/

OpenStudy (anonymous):

What is the value of |x| + y when x = –3 and y = 15? A. –18 B. –12 C. 12 D. 18 would this be C?

OpenStudy (anonymous):

Remember that absolute values are positive values. So, the absolute value of -3 is what? Add what |x| is to y.

OpenStudy (anonymous):

it would be +3 so then 3+15=18? so C?

OpenStudy (anonymous):

Yeppers

OpenStudy (anonymous):

Well no, it's d

OpenStudy (anonymous):

thank you so much again!

OpenStudy (anonymous):

lol ya sorry typed wrong letter.

OpenStudy (anonymous):

Acually this would be C

OpenStudy (anonymous):

oh wait nevermind read the question wrong

OpenStudy (anonymous):

are any of ya'll good at adding and subtracting fractions?